OpenStudy (anonymous):

7(3x + 16) + 8 = -90 what is x?

OpenStudy (anonymous):

the awnser is 10

OpenStudy (anonymous):

7(3x + 16) + 8 = -90 21x+112+8=-90 21x+120=-90 21x=-90-120 21x=-210 x=-210/21 x=-10

OpenStudy (anonymous):

Its not 10 @Moonraven :)

OpenStudy (uri):

@Moonraven Don't give direct answers.